# Quantum field theory treatment of neutrino spin-flavor oscillations in a magnetic field

### v1

@@ 4 Propagators of massive Majorana neutrinos in a magnetic field
Analogously to Eq. (4.5), the rest of the propagators, resulting from Eq. (4), can be derived. We present them for ultrarelativistic neutrinos,
$\Sigma_{11}$ $=\frac{\mathrm{i}(E-E_{2})[1-(\bm{\sigma}\hat{q})]}{2[(E-E_{1}+\mathrm{i}0)(E-E_{2}+\mathrm{i}0)-\mu^{2}B^{2}+\mu^{2}(\hat{q}\mathbf{B})^{2}]},$
$\Sigma_{22}$ $=\frac{\mathrm{i}(E-E_{1})[1-(\bm{\sigma}\hat{q})]}{2[(E-E_{1}+\mathrm{i}0)(E-E_{2}+\mathrm{i}0)-\mu^{2}B^{2}+\mu^{2}(\hat{q}\mathbf{B})^{2}]},$ (4.6)
and $\Sigma_{21}=-\Sigma_{12}$.
5 Transition probability of spin-flavor oscillations
In this section, we calculate the 3D Fourier transform of the dressed propagators derived in Sec. 4, which the matrix element depends on; cf. Eq. (3.11). Then, we compute the transition probability of $\nu_{e}\to\bar{\nu}_{\mu}$ oscillations.
We present the detailed calculation of the $\Sigma_{12}$ contribution to the matrix element. Following Ref. [19], we use the cylindrical coordinates $(\rho,\phi,z)$ to compute the integral in Eq. (3.11), $\mathbf{q}=\rho(\cos\phi\mathbf{e}_{x}+\sin\phi\mathbf{e}_{y})+z\mathbf{e}_{z}$. The spinors of the charged leptons are $\kappa_{e-}^{\mathrm{T}}=(0,1)$ and $\kappa_{\mu+}^{\mathrm{T}}=(1,0)$. Dropping the terms which vanish in the integration over $\phi$, as well as averaging over the leptons states, one gets that
$\kappa_{\mu+}^{\dagger}\Sigma_{12}\kappa_{e-}=$ $\frac{f}{[\mathcal{E}_{+}(\mathbf{q})-\mathcal{E}_{-}(\mathbf{q})]}\left(1+\frac{z}{\sqrt{\rho^{2}+z^{2}}}-\frac{\rho^{2}\cos^{2}\phi}{\rho^{2}+z^{2}}\right)$
×$\times\left(\frac{1}{E-\mathcal{E}_{+}(\mathbf{q})+\mathrm{i}0}-\frac{1}{E-\mathcal{E}_{-}(\mathbf{q})+\mathrm{i}0}\right),$ (5.1)
@@ 5 Transition probability of spin-flavor oscillations
±$\mathcal{E}_{\pm}(\mathbf{q})=\bar{E}\pm\sqrt{\Delta E^{2}+(\mu B)^{2}\left(1-\frac{\rho^{2}\cos^{2}\phi}{\rho^{2}+z^{2}}\right)}.$ (5.3)
@@
Let us compute one of the integrals, which contributes to Eq. (5). Namely
@@
[…]
@@
where $\Delta m^{2}=m_{2}^{2}-m_{1}^{2}>0$. Note that $\mathcal{E}_{+}(E)\neq\mathcal{E}_{+}(\mathbf{q})$; cf. Eq. (5.3). We can approximately set that $\left.\frac{\mathrm{d}\mathcal{E}_{+}(\mathbf{q})}{\mathrm{d}z}\right|{}_{z=z_{+}}\approx\frac{z_{+}}{E}$ and $\mathcal{E}_{+}^{2}(E)\approx\rho^{2}+z_{+}^{2}$. Moreover, for ultrarelativistic leptons, as well as neutrinos, having $E\gg\mathfrak{E}$ and a rather weak magnetic field, obeying $(\mu B)^{2}\ll E\mathfrak{E}$, one has that
$\mathcal{E}_{+}^{2}(E)\approx[\mathcal{E}_{+}^{(0)}(E)]^{2}-\frac{(\mu B)^{2}}{E\mathfrak{E}}\rho^{2}\cos^{2}\phi,$ (5.6)
@@
$\mathcal{E}_{+}^{(0)}(E)=E+\mathfrak{E},\quad\mathfrak{E}=\sqrt{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}}.$ (5.7)
Using Eq. (5.6), one gets that the poles $z_{+}$ obey the relation,
$[\mathcal{E}_{+}^{(0)}(E)]^{2}=\rho^{2}\left(1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi\right)+z_{+}^{2}.$ (5.8)
Finally, we obtain the following expressions for the poles in the explicit form:
@@
$I_{+}=$ $\frac{\mathrm{i}\mu BE}{16\pi^{2}\mathfrak{E}}\int_{0}^{2\pi}\frac{\mathrm{d}\phi}{\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}}\Bigg{[}\int_{0}^{\rho_{0}}\frac{\rho\mathrm{d}\rho\exp\left(\mathrm{i}\tilde{L}\sqrt{\rho_{0}^{2}-\rho^{2}}\right)}{\sqrt{\rho_{0}^{2}-\rho^{2}}}$
×$\times\left(1+\frac{\sqrt{\rho_{0}^{2}-\rho^{2}}}{\tilde{E}}-\frac{\rho^{2}\cos^{2}\phi}{E^{2}}\left[1-\frac{(\mu B)^{2}}{2\mathfrak{E}^{2}}\right]\right)$
$-\mathrm{i}\int_{\rho_{0}}^{\infty}\frac{\rho\mathrm{d}\rho\exp\left(-\tilde{L}\sqrt{\rho^{2}-\rho_{0}^{2}}\right)}{\sqrt{\rho^{2}-\rho_{0}^{2}}}\left(1+\mathrm{i}\frac{\sqrt{\rho^{2}-\rho_{0}^{2}}}{\tilde{E}}-\frac{\rho^{2}\cos^{2}\phi}{E^{2}}\left[1-\frac{(\mu B)^{2}}{2\mathfrak{E}^{2}}\right]\right)\Bigg{]},$ (5.10)
@@
$I_{+}\approx$ $\frac{\mu BEe^{i(E+\mathfrak{E})}}{8\pi{}^{2}\mathfrak{E}L}\left(1+\frac{\mathfrak{E}}{2E}\right)\int_{0}^{2\pi}\frac{\mathrm{d}\phi}{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}$
$\approx\frac{\mu BEe^{\mathrm{i}(E+\mathfrak{E})L}}{4\pi\mathfrak{E}L}\left(1+\frac{\left(\frac{\Delta m^{2}}{4E}\right)^{2}}{2\mathfrak{E}E}\right),$ (5.12)
@@
$I_{-}\approx\frac{\mu BEe^{\mathrm{i}(E-\mathfrak{E})L}}{4\pi\mathfrak{E}L}\left(1-\frac{\left(\frac{\Delta m^{2}}{4E}\right)^{2}}{2\mathfrak{E}E}\right).$ (5.14)
The contributions of the diagonal propagators $\Sigma_{aa}$ in Eq. (4) to the matrix element $\mathcal{M}_{e\to\bar{\mu}}$ are vanishing.
@@
$\mathcal{M}_{e\to\bar{\mu}}=I_{-}-I_{+}=-\frac{\mathrm{i}\mu BEe^{\mathrm{i}EL}}{2\pi\mathfrak{E}L}\left(\sin\mathfrak{E}L-\frac{\mathrm{i}\left(\Delta m^{2}\right)^{2}}{32\mathfrak{E}E^{3}}\cos\mathfrak{E}L\right).$ (5.15)
The probability of the process $\nu_{e}\to\bar{\nu}_{\mu}$ is $P_{\nu_{e}\to\bar{\nu}_{\mu}}\propto|\mathcal{M}_{e\to\bar{\mu}}|^{2}$. Based on Eq. (5.15) and accounting for Eq. (5.7), we obtain that $P_{\nu_{e}\to\bar{\nu}_{\mu}}=P_{\nu_{e}\to\bar{\nu}_{\mu}}^{(0)}+\delta P_{\nu_{e}\to\bar{\nu}_{\mu}}$, where
$P_{\nu_{e}\to\bar{\nu}_{\mu}}^{(0)}(L)$ $=\frac{(\mu B)^{2}}{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}}\sin^{2}\left(\sqrt{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}}L\right),$ (5.16)
$\delta P_{\nu_{e}\to\bar{\nu}_{\mu}}(L)$ $=\frac{(\mu B)^{2}(\Delta m^{2})^{4}}{4096E^{6}\left[(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}\right]^{2}}\sin^{2}\left(2\sqrt{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}}L\right).$ (5.17)
We compare the transition probability in Eq. (5.17) with the result of the quantum mechanical approach shortly in Sec. 6.
6 Quantum mechanical description of spin-flavor oscillations

### v2

@@ 4 Propagators of massive Majorana neutrinos in a magnetic field
Analogously to Eq. (4.5), one gets that $\Sigma_{21}=-\Sigma_{12}$.
5 Transition probability of the spin-flavor precession
In this section, we calculate the 3D Fourier transform of the dressed propagators, derived in Sec. 4, which the matrix element depends on; cf. Eq. (3.11). Then, we compute the transition probability for the $\nu_{e}\to\bar{\nu}_{\mu}$ precession.
We present the detailed calculation of the $\Sigma_{12}$ contribution to the matrix element. The magnetic field is taken to be transverse with respect $\mathbf{L}$, $\mathbf{B}\perp\mathbf{L}$. For instance, we suppose that $\mathbf{L}=L\mathbf{e}_{z}$ and $\mathbf{B}=B\mathbf{e}_{x}$, where $\mathbf{e}_{x,z}$ are the unit vectors along the $x$- and $z$-axes. In the quantum mechanical approach, a neutrino is on the mass shell, i.e. the neutrino momentum is also perpendicular to $\mathbf{B}$. In the QFT treatment, a neutrino is a virtual particle. Thus, $\mathbf{q}$ is arbitrary and, hence, $(\hat{q}\cdot\mathbf{B})\neq 0$ in Eq. (4.5).
We should keep the terms (i) $\propto\mu(\bm{\sigma}\cdot\hat{q})(\hat{q}\cdot\mathbf{B})$ and (ii) $\propto\mu^{2}(\hat{q}\cdot\mathbf{B})^{2}$ in the numerator and the denominator of $\Sigma_{12}$ in Eq. (4.5) since this kind of terms is of the same order of magnitude as other terms in the propagator. At least, one has no reasons to neglect the terms (i) and (ii) before the calculation of the matrix element.
To evaluate the contribution of the terms (i) and (ii) to the matrix element we adopt the forward scattering approximation, i.e. we assume that $\mathbf{p}_{e}\upuparrows\mathbf{p}_{\bar{\mu}}\upuparrows\mathbf{L}$ to simplify some integrals computation. This approximation is valid for ultrarelativistic charged leptons. The leptons with arbitrary energies are considered shortly in Sec. 5.1.
Following Ref. [20], we use the cylindrical coordinates $(\rho,\phi,z)$ to compute the integral in Eq. (3.11), $\mathbf{q}=\rho(\cos\phi\mathbf{e}_{x}+\sin\phi\mathbf{e}_{y})+z\mathbf{e}_{z}$. In the forward scattering approximation, the spinors of the charged leptons are $\kappa_{e\mathrm{L}}^{\mathrm{T}}=(0,1)$ and $\kappa_{\bar{\mu}\mathrm{R}}^{\mathrm{T}}=(1,0)$. After averaging the contribution of $\Sigma_{12}$ to the matrix element over $\kappa_{\bar{\mu}\mathrm{R}}$ and $\kappa_{e\mathrm{L}}$, we can drop some terms which vanish in the integration over $\phi$ because of the symmetry properties of the integrand. Therefore, the remainung nontrivial contribution reads
$\kappa_{\bar{\mu}\mathrm{R}}^{\dagger}\Sigma_{12}\kappa_{e\mathrm{L}}=\frac{f}{\mathcal{E}_{+}(\mathbf{q})-\mathcal{E}_{-}(\mathbf{q})}\left(\frac{1}{E-\mathcal{E}_{+}(\mathbf{q})+\mathrm{i}0}-\frac{1}{E-\mathcal{E}_{-}(\mathbf{q})+\mathrm{i}0}\right),$ (5.1)
@@ 5 Transition probability of the spin-flavor precession
$\mathcal{E}_{\pm}(\mathbf{q})=\bar{E}\pm\sqrt{\Delta E^{2}+(\mu B)^{2}\left(1-\frac{\rho^{2}\cos^{2}\phi}{\rho^{2}+z^{2}}\right)}.$ (5.3)
@@
Let us compute one of the integrals, which contributes to Eq. (5.1). Namely
@@
where $z_{+}(\rho,\phi)$ are the roots of the equation $E-\mathcal{E}_{+}(\rho,\phi,z)+\mathrm{i}0=0$. These roots are in the upper half-plane. Since we take that $L>0$, we close the contour in the upper half-plane while integrating over $z$.
@@
where $\Delta m^{2}=m_{2}^{2}-m_{1}^{2}>0$. Note that $\mathcal{E}_{+}(E)\neq\mathcal{E}_{+}(\mathbf{q})$; cf. Eq. (5.3). We can approximately set that $\left.\frac{\mathrm{d}\mathcal{E}_{+}(\mathbf{q})}{\mathrm{d}z}\right|{}_{z=z_{+}}\approx\frac{z_{+}}{\mathcal{E}_{+}(E)}$ and $\mathcal{E}_{+}^{2}(E)\approx\rho^{2}+z_{+}^{2}$. Moreover, for ultrarelativistic neutrinos, having $\mathfrak{E}\ll E$, and a rather weak magnetic field, obeying $(\mu B)^{2}\ll E\mathfrak{E}$, one has that
$\mathcal{E}_{+}(E)\approx E+\mathfrak{E}\left(1-\frac{(\mu B)^{2}\rho^{2}\cos^{2}\phi}{2E^{2}\mathfrak{E}^{2}}\right),$ (5.6)
@@
$\mathfrak{E}=\sqrt{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}}.$ (5.7)
Using Eq. (5.6), one gets the following expressions for the poles in the explicit form:
@@
$I_{+}=$ $\frac{\mathrm{i}\mu B(E+\mathfrak{E})}{16\pi^{2}\mathfrak{E}}\int_{0}^{2\pi}\frac{\mathrm{d}\phi}{\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}}\Bigg{[}\int_{0}^{\rho_{0}}\frac{\rho\mathrm{d}\rho\exp\left(\mathrm{i}\tilde{L}\sqrt{\rho_{0}^{2}-\rho^{2}}\right)}{\sqrt{\rho_{0}^{2}-\rho^{2}}}$
$\times\left(1-\rho^{2}\cos^{2}\phi\left(\frac{1}{(E+\mathfrak{E})^{2}}-\frac{(\mu B)^{2}}{2E^{2}\mathfrak{E}^{2}}\right)+\frac{\sqrt{\rho_{0}^{2}-\rho^{2}}}{\tilde{E}\left(1+\frac{\mathfrak{E}}{E}\right)}\right)$
$-\mathrm{i}\int_{\rho_{0}}^{\infty}\frac{\rho\mathrm{d}\rho\exp\left(-\tilde{L}\sqrt{\rho^{2}-\rho_{0}^{2}}\right)}{\sqrt{\rho^{2}-\rho_{0}^{2}}}$
$\times\left(1-\rho^{2}\cos^{2}\phi\left(\frac{1}{(E+\mathfrak{E})^{2}}-\frac{(\mu B)^{2}}{2E^{2}\mathfrak{E}^{2}}\right)+\frac{\mathrm{i}\sqrt{\rho^{2}-\rho_{0}^{2}}}{\tilde{E}\left(1+\frac{\mathfrak{E}}{E}\right)}\right)\Bigg{]},$ (5.9)
@@
$I_{+}\approx$ $\frac{\mu B(E+\mathfrak{E})e^{\mathrm{i}(E+\mathfrak{E})}}{8\pi{}^{2}\mathfrak{E}L}\int_{0}^{2\pi}\frac{\mathrm{d}\phi}{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}$
$\approx\frac{\mu B(E+\mathfrak{E})e^{\mathrm{i}(E+\mathfrak{E})L}}{4\pi\mathfrak{E}L}\left(1-\frac{(\mu B)^{2}}{2E\mathfrak{E}}\right),$ (5.11)
@@
$I_{-}\approx\frac{\mu B(E-\mathfrak{E})e^{\mathrm{i}(E-\mathfrak{E})L}}{4\pi\mathfrak{E}L}\left(1-\frac{(\mu B)^{2}}{2E\mathfrak{E}}\right).$ (5.13)
As we mentioned in Sec. 4, the contributions of the diagonal propagators $\Sigma_{aa}$ to the matrix element $\mathcal{M}_{e\to\bar{\mu}}$ are vanishing.
@@
$\mathcal{M}_{e\to\bar{\mu}}=I_{-}-I_{+}=-\frac{\mathrm{i}Ee^{\mathrm{i}EL}}{2\pi L}\frac{\mu B}{\mathfrak{E}}\left(1-\frac{(\mu B)^{2}}{2\mathfrak{E}E}\right)\left(\sin\mathfrak{E}L-\mathrm{i}\frac{\mathfrak{E}}{E}\cos\mathfrak{E}L\right).$ (5.14)
The probability of the process $\nu_{e}\to\bar{\nu}_{\mu}$ is $P_{\nu_{e}\to\bar{\nu}_{\mu}}\propto|\mathcal{M}_{e\to\bar{\mu}}|^{2}$. Based on Eq. (5.14), we obtain that
$P_{\nu_{e}\to\bar{\nu}_{\mu}}(L)\approx P_{\mathrm{max}}\sin^{2}\left(\sqrt{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}}L\right),$ (5.15)
where the amplitude of the transition probability has two terms, $P_{\mathrm{max}}=P_{\mathrm{max}}^{(\mathrm{qm})}+P_{\mathrm{max}}^{(\mathrm{qft})}$. These terms have the following forms:
$P_{\mathrm{max}}^{(\mathrm{qm})}=\frac{(\mu B)^{2}}{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}},\quad P_{\mathrm{max}}^{(\mathrm{qft})}=-\frac{(\mu B)^{4}}{E\left[(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}\right]^{3/2}}.$ (5.16)
We shall see shortly in Sec. 6 that $P_{\mathrm{max}}^{(\mathrm{qm})}$ in Eq. (5.16) reproduces the result of the quantum mechanical approach. The term $P_{\mathrm{max}}^{(\mathrm{qft})}$ in Eq. (5.16) is the QFT contribution to the transition probability. It arises from the terms (i) and (ii) in the propagator $\Sigma_{12}$
To derive Eqs. (5.15) and (5.16), we take into account Eq. (5.7) and assume that $\mu B\sim\mathfrak{E}$. Moreover, we neglect the terms $\sim(\mathfrak{E}/E)^{2}\ll 1$, whereas we keep the terms $\sim\mu B/E$ or $\sim\mathfrak{E}/E$.
One can see in Eq. (5.16) that the QFT contribution to the transition probability is negative. Moreover, $|P_{\mathrm{max}}^{(\mathrm{qft})}|/P_{\mathrm{max}}^{(\mathrm{qm})}\sim\mu B/E\sim\mathfrak{E}/E$. Thus, the QFT term is much smaller than the quantum mechanical one for all realistic oscillations channels where neutrinos are ultrarelativistic.
The latter result proves the assumption that the terms (i) and (ii) in $\Sigma_{12}$ in Eq. (4.5) can be safely neglected for ultrarelativistic neutrinos since their contribution to the transition probability is small. We recall that all the terms in $\Sigma_{12}$ in Eq. (4.5), including (i) and (ii), are of the same order of magnitude. Therefore, before the transition probability computation, one cannot say whether certain terms in $\Sigma_{12}$ are smaller or greater.
5.1 Scattering of charged leptons with arbitrary energies
We have just demonstrated that the contributions of the terms (i) and (ii) in the propagator in Eq. (4.5), which have the QFT origin, to the transition probability of the spin-flavor precession are small. The washing out of these terms happens for ultrarelativistic neutrinos. That is why, we can neglect the terms (i) and (ii) while we study the case of charged leptons with arbitrary energies, with neutrinos being ultrarelativistic.
For simplicity, we take that an incoming electron still propagates along $\mathbf{L}$. The outgoing antimuon can have an arbitrary direction of its momentum. The spin of an electron is taken to be opposite to the electron momentum, $\mathbf{s}_{e}\downuparrows\mathbf{p}_{e}$, whereas the antimuon spin is supposed to be along the particle momentum, $\mathbf{s}_{\bar{\mu}}\upuparrows\mathbf{p}_{\bar{\mu}}$.
The chiral projections of leptons bispinors have the forms,
$\kappa_{e\mathrm{L}}=\sqrt{\frac{1+v_{e}}{2}}\left(\begin{array}[]{c}0\\ 1\end{array}\right),\quad\kappa_{\bar{\mu}\mathrm{R}}=-\sqrt{\frac{1+v_{\bar{\mu}}}{2}}\left(\begin{array}[]{c}e^{-\mathrm{i}\phi_{\bar{\mu}}}\cos\frac{\theta_{\bar{\mu}}}{2}\\ e^{\mathrm{i}\phi_{\bar{\mu}}}\sin\frac{\theta_{\bar{\mu}}}{2}\end{array}\right),$ (5.17)
where $v_{e}$ and $v_{\bar{\mu}}$ are the velocities of the electron and the antimuon, and the spherical angles $\phi_{\bar{\mu}}$ and $\theta_{\bar{\mu}}$ fix the direction of the antimuon momentum with respect to $\mathbf{e}_{z}$.
The averaging of $\Sigma_{12}$ over $\kappa_{\bar{\mu}\mathrm{R}}$ and $\kappa_{e\mathrm{L}}$ is similar to that in Eq. (5.1) except the following replacements:
$f(\rho,\phi,z)\to$ $\sqrt{(1+v_{e})(1+v_{\bar{\mu}})}e^{\mathrm{i}\phi_{\bar{\mu}}}\cos\left(\frac{\theta_{\bar{\mu}}}{2}\right)\frac{\mu B}{4}\left(1+\frac{z}{\sqrt{\rho^{2}+z^{2}}}\right),$
$\mathcal{E}_{\pm}(\mathbf{q})\to$ $\bar{E}\pm\sqrt{\Delta E^{2}+(\mu B)^{2}}.$ (5.18)
The subsequent computation of the matrix element is identical to that in Sec. 5. The final result reads
@@ 5.1 Scattering of charged leptons with arbitrary energies
$\mathcal{M}_{e\to\bar{\mu}}=\frac{\mathrm{i}Ee^{\mathrm{i}EL}}{4\pi L}\sqrt{(1+v_{e})(1+v_{\bar{\mu}})}e^{\mathrm{i}\phi_{\bar{\mu}}}\cos\left(\frac{\theta_{\bar{\mu}}}{2}\right)\frac{\mu B}{\mathfrak{E}}\sin\mathfrak{E}L,$ (5.19)
i.e. we remove the QFT corrections from Eq. (5.14) and take into account the additional factor $e^{\mathrm{i}\phi_{\bar{\mu}}}\cos\left(\tfrac{\theta_{\bar{\mu}}}{2}\right)\sqrt{(1+v_{e})(1+v_{\bar{\mu}})}/2$.
The measurable quantity of the process $e+N\to\bar{\mu}+\tilde{N}^{\prime}$ is the differential cross-section [25, p. 252],
$\mathrm{d}\sigma_{e\to\bar{\mu}}=G_{\mathrm{int}}^{4}E^{2}|\mathcal{M}_{e\to\bar{\mu}}|^{2}\frac{v_{\bar{\mu}}}{v_{e}}\mathrm{d}\Omega_{\bar{\mu}},$ (5.20)
where $\mathrm{d}\Omega_{\bar{\mu}}=2\pi\sin\theta_{\bar{\mu}}\mathrm{d}\theta_{\bar{\mu}}$ is the solid angle of the antimuon momentum space.
Integrating Eq. (5.20) over $\mathrm{d}\Omega_{\bar{\mu}}$, one gets that the total cross-section is
$\sigma_{e\to\bar{\mu}}=\int\mathrm{d}\sigma_{e\to\bar{\mu}}=\frac{E^{4}G_{\mathrm{int}}^{4}}{2\pi L^{2}}P^{(\mathrm{eff})}_{\nu_{e}\to\bar{\nu}_{\mu}},\quad P^{(\mathrm{eff})}_{\nu_{e}\to\bar{\nu}_{\mu}}=(1+v_{e})(1+v_{\bar{\mu}})\frac{v_{\bar{\mu}}}{4v_{e}}\left(\frac{\mu B}{\mathfrak{E}}\right)^{2}\sin^{2}(\mathfrak{E}L).$ (5.21)
One can see in Eq. (5.21) that the total cross-section oscillates in space and is proportional to the quantum mechanical transition probability, $P_{\mathrm{max}}^{(\mathrm{qm})}\sin^{2}(\mathfrak{E}L)$; cf. Eqs. (5.15) and (5.16).
For ultrarelativistic charged leptons, $P^{(\mathrm{eff})}_{\nu_{e}\to\bar{\nu}_{\mu}}\approx P^{(\mathrm{qm})}_{\nu_{e}\to\bar{\nu}_{\mu}}$ since $v_{\bar{\mu}}\sim v_{e}\sim 1$. In general situation, $P^{(\mathrm{eff})}_{\nu_{e}\to\bar{\nu}_{\mu}}<P^{(\mathrm{qm})}_{\nu_{e}\to\bar{\nu}_{\mu}}$ since $\mathfrak{m}_{\mu}>\mathfrak{m}_{e}$. Thus, the factor $(1+v_{e})(1+v_{\bar{\mu}})v_{\bar{\mu}}/(4v_{e})<1$ should be accounted for while designing a neutrino oscillations experiment.
6 Quantum mechanical description of the spin-flavor precession